# A Library of LLM Intrinsics for Retrieval-Augmented Generation

### v1

@@ Abstract
In the developer community for large language models (LLMs), there is not yet a clean pattern analogous to a software library, to support very large scale collaboration. Even for the commonplace use case of Retrieval-Augmented Generation (RAG), it is not currently possible to write a RAG application against a well-defined set of APIs that are agreed upon by different LLM providers. Inspired by the idea of compiler intrinsics, we propose some elements of such a concept through introducing a library of LLM Intrinsics for RAG. An LLM intrinsic is defined as a capability that can be invoked through a well-defined API that is reasonably stable and independent of how the LLM intrinsic itself is implemented. The intrinsics in our library are released as LoRA adapters on HuggingFace, and through a software interface with clear structured input/output characteristics on top of vLLM as an inference platform, accompanied in both places with documentation and code. This article describes the intended usage, training details, and evaluations for each intrinsic, as well as compositions of multiple intrinsics.
@@ 1.1 Overview of the RAG LLM Intrinsics Library
The LLM Intrinsics RAG Library currently comprises five intrinsics, each of which expects as input a (single-turn or multi-turn) conversation between a user and an AI assistant. Three of the intrinsics also expect a set of grounding passages. The functionality of each intrinsic is described below, and Table 1 summarizes the inputs and outputs of each one.
Query Rewrite (QR). Given a conversation ending with a user query, QR will decontextualize that last user query by rewriting it (whenever necessary) into an equivalent version that is standalone and can be understood by itself. While this adapter is general purpose for any multi-turn conversation, it is especially effective in RAG settings where its ability to rewrite a user query into a standalone version directly improves the retriever performance, which in turn improves the answer generation performance. This is a pre-generation intrinsic since its suggested use is before invoking retrieval.
@@
Hallucination Detection (HD). Given a conversation ending with an assistant response, and a set of passages, HD outputs a hallucination risk range for each sentence in the last assistant response, with respect to the set of passages. It could be used in concert with sampling techniques that yield multiple generated responses, some of which could then be filtered according to their HD scores. This is a post-generation intrinsic since its expected use is after invoking the LLM to create the response.
@@
Intrinsic Input Output Pre/Post Gen
Passages Query Resp.
Query Rewrite (QR) $\times$ Standalone version of last query pre
Uncertainty Quantification (UQ) $\times$ (optional) $\times$ Certainty score for last assistant response (before generation) pre
$\times$ (optional) $\times$ Certainty score for last assistant response (after generation) post
Hallucination Detection (HD) $\times$ $\times$ Hallucination score for last assistant response post
Answerability Determination (AD) $\times$ $\times$ Flag denoting if last query is answerable from passages pre
Citation Generation (CG) $\times$ $\times$ Citations for last assistant response based on passages post
Table 1: RAG LLM Intrinsics with their expected inputs and outputs. Query and Resp. refer to conversations ending with a user query and assistant response, respectively. Pre/Post Gen denotes if an intrinsic is called before or after generation, respectively.
@@ 1.2 RAG LLM Intrinsics Implementation
Each of these intrinsics has been implemented by training a LoRA adapter for ibm-granite/granite-3.2-8b-instruction fine-tuned for a particular task. Each of these LoRA models has been released on HuggingFace:
@@ 2 Query Rewrite
Granite 3.2 8b Instruct - Query Rewrite is a LoRA adapter for ibm-granite/granite-3.2-8b-instruct fine-tuned for the following task:
@@ 2.1 Intended Use
See Section A.1 for an example describing how to use the Query Rewrite intrinsic.
@@ 2.2.1 Evaluation of the retriever
Granite 3.2-8b-instruct-query-rewrite-LoRA 0.56 0.68 0.76
@@
Granite 3.2-8b-instruct-query-rewrite-LoRA 0.44 0.57 0.66
@@
Granite 3.2-8b-instruct-query-rewrite-LoRA 0.63 0.75 0.83
@@
If we focus on Recall@20 numbers, as one instance of the metric, there is an overall 9 percentage points jump when using query rewrite with the Granite 3.2-8b LoRA adapter versus when using the no rewrite strategy. This jump is more pronounced on the non-standalone fragment, where query rewrite with the Granite 3.2-8b LoRA adapter leads to 22 percentage points improvement over the no-rewrite strategy. Also, we can observe that the numbers with the LoRA rewrites are very close to what can be obtained with the gold rewrites on non-standalones (and slightly better on standalones for LoRA – human annotators were instructed to leave the query unchanged when classifying it as standalone, however, the LoRA adapter may still perform some rewriting which turns out to further improve the recall).
@@ 2.2.2 Evaluation of answer generation
We evaluate answer generation quality, with top-k passages retrieved under the various query rewrite strategies for the retriever. We choose here k = 20, but similar trends take place for other values of k. We used Granite-3.2-8b instruct as the answer generator, and RAGAS Faithfulness (RAGAS-F) and RAD-Bench score as metrics for answer quality. We use the same three testsets as above.
@@
Granite 3.2-8b-instruct-query-rewrite-LoRA 0.81 0.70
@@
Granite 3.2-8b-instruct-query-rewrite-LoRA 0.83 0.71
@@
As with Recall, similar observations can be made here as well. Specifically, we see an 8 percentage points jump in RAGAS Faithfulness and 4 percentage points jump in RAD-Bench score when using query rewrite with the Granite 3.2-8b LoRA adapter versus when using the no rewrite strategy. This improvement is more pronounced on the non-standalone fragment, where query rewrite with the Granite 3.2-8b LoRA adapter leads to a 18 percentage points jump in RAGAS Faithfulness and 7 percentage points jump in RAD-Bench score.

### v2

@@ Abstract
In the developer community for large language models (LLMs), there is not yet a clean pattern analogous to a software library, to support very large scale collaboration. Even for the commonplace use case of Retrieval-Augmented Generation (RAG), it is not currently possible to write a RAG application against a well-defined set of APIs that are agreed upon by different LLM providers. Inspired by the idea of compiler intrinsics, we propose some elements of such a concept through introducing a library of LLM Intrinsics for RAG. An LLM intrinsic is defined as a capability that can be invoked through a well-defined API that is reasonably stable and independent of how the LLM intrinsic itself is implemented. The intrinsics in our library are released as LoRA adapters on HuggingFace, and through a software interface with clear structured input/output characteristics on top of vLLM as an inference platform, accompanied in both places with documentation and code. This article describes the intended usage, training details, and evaluations for each intrinsic, as well as compositions of multiple intrinsics. [^1] This (June 2025) is the second version of this paper (the first was published in April 2025). Intrinsics implemented as LoRAs are now trained on IBM Granite 3.3 8b instruct (previously 3.2)
@@ 1.1 Overview of the RAG LLM Intrinsics Library
The LLM Intrinsics RAG Library currently comprises eight intrinsics, each of which expects as input a (single-turn or multi-turn) conversation between a user and an AI assistant. Six of the intrinsics also expect a set of grounding passages. The functionality of each intrinsic is described below, and Table 1 summarizes the inputs and outputs of each one.
Query Rewrite (QR). Given a conversation ending with a user query, QR will decontextualize that last user query by rewriting it (whenever necessary) into an equivalent version that is standalone and can be understood by itself. While this adapter is general purpose for any multi-turn conversation, it is especially effective in RAG settings where its ability to rewrite a user query into a standalone version directly improves the retriever performance, which in turn improves the answer generation performance. This is a pre-retrieval intrinsic since its suggested use is before invoking retrieval.
Query Expansion (QE). Given a conversation ending with a user query, QE is designed to probe the retriever from multiple angles by generating a set of semantically diverse versions of that last user query. This expanded set of queries provides diverse retrieval paths, and thus this intrinsic is particularly effective in RAG settings, especially with terse, general, or underspecified queries. Like Query Rewrite, this is a pre-retrieval intrinsic.
Context Relevance (CR). Given a conversation ending with a user query, and an individual passage, CR classifies whether the passage is relevant, partially relevant, or irrelevant for answering the last user query - or if the passage may instead mislead or harm the downstream generator model’s response quality. This is a pre-generation intrinsic.
Passage Reranking (PRR). Given a conversation ending with a user query, and a set of passages, PRR returns a ranked list of the passages ordered by suitability to answering the query. If the number of passages is small ($<10$) all passages are compared pairwise and returned ranked by win count; otherwise a tournament algorithm is used. This is a pre-generation intrinsic.
@@
Hallucination Detection (HD). Given a conversation ending with an assistant response, and a set of passages, HD outputs a hallucination risk for each sentence in the last assistant response, with respect to the set of passages. It could be used in concert with sampling techniques that yield multiple generated responses, some of which could then be filtered according to their hallucination risks. This is a post-generation intrinsic since its expected use is after invoking the LLM to create the response.
@@
Input
Intrinsic Passages End Query End Resp. Output Pre/Post
Query Rewrite (QR) $\times$ Standalone version of last query Pre-R
Query Expansion (QE) $\times$ Array of query variants Pre-R
Context Relevance (CR) $\times$ (1 passage) $\times$ Flag denoting if passage is relevant to query Pre-G
Answerability Determination (AD) $\times$ $\times$ Flag denoting if last query is answerable from passages Pre-G
Passage Reranking (PRR) $\times$ $\times$ Reranked list of passages Pre-G
Uncertainty Quantification (UQ) $\times$ (optional) $\times$ Certainty score for last assistant response (before generation) Pre-G
$\times$ (optional) $\times$ Certainty score for last assistant response (after generation) Post-G
Hallucination Detection (HD) $\times$ $\times$ Hallucination risk for last assistant response Post-G
Citation Generation (CG) $\times$ $\times$ Citations for last assistant response based on passages Post-G
Table 1: RAG LLM Intrinsics with their expected inputs and outputs. End Query and End Resp. refer to conversations ending with a user query and assistant response, respectively. Pre/Post denotes if an intrinsic is called before retrieval (Pre-R), after retrieval and before generation (Pre-G), or after generation (Post-G).
@@ 1.2 RAG LLM Intrinsics Implementation
Six of these intrinsics have been implemented by training a LoRA adapter for ibm-granite/granite-3.3-8b-instruct fine-tuned for a particular task. These have been released on HuggingFace as the Granite 3.3 8b RAG Agent Library, available at: https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib. The six intrinsics thus implemented are Query Rewrite, Context Relevance, Answerability Determination, Uncertainty Quantification, Hallucination Detection, and Citation Generation.
@@
Furthermore, the remaining two intrinsics, Query Expansion and Passage Reranking are pure software-based implementations and are only available through Granite IO. This emphasizes the general nature of intrinsics, where the implementation is abstracted from the function, as previously discussed.
@@ 2 Query Rewrite
Granite 3.3 8b Instruct - Query Rewrite is a LoRA adapter for ibm-granite/granite-3.3-8b-instruct fine-tuned for the following task:
@@ 2.1 Intended Use
For an example describing how to use the intrinsic, refer to the documentation of the RAG agent library at https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib.
@@ 2.2.1 Evaluation of the retriever
Granite 3.3-8b-instruct-query-rewrite-LoRA 0.56 0.68 0.76
@@
Granite 3.3-8b-instruct-query-rewrite-LoRA 0.44 0.57 0.66
@@
Granite 3.3-8b-instruct-query-rewrite-LoRA 0.63 0.75 0.83
@@
If we focus on Recall@20 numbers, as one instance of the metric, there is an overall 9 percentage points jump when using query rewrite with the Granite 3.3-8b LoRA adapter versus when using the no rewrite strategy. This jump is more pronounced on the non-standalone fragment, where query rewrite with the Granite 3.3-8b LoRA adapter leads to 22 percentage points improvement over the no-rewrite strategy. Also, we can observe that the numbers with the LoRA rewrites are very close to what can be obtained with the gold rewrites on non-standalones (and slightly better on standalones for LoRA – human annotators were instructed to leave the query unchanged when classifying it as standalone, however, the LoRA adapter may still perform some rewriting which turns out to further improve the recall).
@@ 2.2.2 Evaluation of answer generation
We evaluate answer generation quality, with top-k passages retrieved under the various query rewrite strategies for the retriever. We choose here k = 20, but similar trends take place for other values of k. We used Granite-3.3-8b instruct as the answer generator, and RAGAS Faithfulness (RAGAS-F) and RAD-Bench score as metrics for answer quality. We use the same three testsets as above.
@@
Granite 3.3-8b-instruct-query-rewrite-LoRA 0.81 0.70
@@
Granite 3.3-8b-instruct-query-rewrite-LoRA 0.83 0.71
@@
As with Recall, similar observations can be made here as well. Specifically, we see an 8 percentage points jump in RAGAS Faithfulness and 4 percentage points jump in RAD-Bench score when using query rewrite with the Granite 3.3-8b LoRA adapter versus when using the no rewrite strategy. This improvement is more pronounced on the non-standalone fragment, where query rewrite with the Granite 3.3-8b LoRA adapter leads to a 18 percentage points jump in RAGAS Faithfulness and 7 percentage points jump in RAD-Bench score.
@@ 2.3 Training Details
3 Query Expansion
While Query Rewrite transforms a conversational query into a self-contained standalone form using a fine-tuned LoRA adapter, we introduce a Query Expansion intrinsic that takes this further by generating a set of semantically diverse queries designed to probe the retriever from multiple angles. Instead of relying on a single rewrite, this intrinsic generates multiple candidate queries. These reflect different interpretations or formulations of the original user intent, improving the likelihood of retrieving relevant supporting passages. Given a multi-turn conversation, this intrinsic outputs a list of expanded queries by employing several expansion strategies:
  • last user-turn question (no rewrites)
  • Query Rewrite LoRA
  • Answer Sampling: Samples a plausible answer using the granite instruct model without using any documents.
  • Backward Generation from Answer: Reverse-engineer a potential question that could have elicited the answer.
  • Synonymic Rewrite: Prompts Granite with explicit instructions to generate synonymous version of the query.
3.1 INTENDED USE
The query expansion intrinsic is designed for multi-turn conversational retrieval settings, particularly those involving RAG pipelines. Its intended use is to generate multiple semantically varied formulations of the final user query, enabling more diverse retrieval. The expanded queries can be used independently to query a retriever. The expanded queries provide diverse retrieval paths, improving the chances of surfacing relevant supporting content from the documents.
Input: The input to the model is a list of conversational turns. These turns can alternate between the user and assistant roles, and the last turn is assumed to be from the user.
Output: List of expanded queries (with user role).
For an example describing how to use the intrinsic, refer to the relevant documentation in Granite IO at https://github.com/ibm-granite/granite-io/tree/main/notebooks.
3.2 EVALUATION
3.2.1 EVALUATION OF THE RETRIEVER
We evaluate retrieval performance using Recall@k on the MT-RAG benchmark, comparing Query Expansion against a range of querying baselines in Table 8. The retrieval backend is the Elser retriever, kept fixed across all experiments to isolate the impact of query formulation. With query expansion method, each expanded query is used to retrieve top-$20$ passages using Elser retriever. The final set of retrieved passages is obtained by taking set union across all querying methods, yielding a maximum of $100$ passages.
Method Recall@50 Recall@100
Last user-turn (No rewrite) 74 77
Query Rewrite LoRA 84 87
Answer-1 78 85
Reverse-Engineered Question 79 84
Synonymous Query Rewrite 62 69
Query Expansion 87 89
Table 8: Comparison of different querying strategies on the retrieval task on full MT-RAG dataset
We observe that Query Expansion outperforms all single-query baselines, achieving Recall of 87% with top-$50$ retrieved passages, compared to 84% for the best-performing individual method (Query Rewrite LoRA). At Recall@100, the margin remains consistent, with Query Expansion reaching 89%, showing that expanded queries retrieve relevant documents earlier. Beyond improvements in Recall@k, we observe distinct qualitative benefits of Query Expansion over traditional single-query methods: examples where the user’s last-turn query is vague or underspecified. In such cases, a single rewrite often commits to one interpretation, which may miss relevant documents that align with alternate plausible meanings. By contrast, Query Expansion generates multiple semantically distinct versions of the query, each capturing a different potential reading of the user’s intent.
Method RAGAS-F@50 RAGAS-F@100 RAD-Bench@50 RAD-Bench@100
Last user-turn (No rewrite) 77 82 66 66
Query Rewrite LoRA 83 86 68 67
Answer-1 85 86 69 68
Reverse-Engineered Question 82 86 67 67
Synonymous Query Rewrite 75 80 67 66
Query Expansion 85 85 67 67
Table 9: Comparison of different querying strategies on the answer quality on full MT-RAG dataset
3.2.2 EVALUATION OF ANSWER GENERATION
We evaluate answer generation quality, with top-k passages retrieved under the various querying strategies for the retriever. We used Granite3.3-8b instruct as the answer generator, and RAGAS Faithfulness (RAGAS-F) and RAD-Bench score as metrics for answer quality. Table 9 shows the results. We find that methods like Answer-1 and Query Rewrite LoRA yield better faithfulness and correctness scores, indicating synergy in retrieval and generation scores. Last user-turn (No rewrite) shows the lowest generation quality scores related to other methods, highlighting the benefits of query modification strategies. Query Expansion maintains competitive performance. Importantly, it does not cause any regression in answer quality relative to other methods.
Recommended usecases.
  • This can be especially useful for ambiguous or underspecified queries, where a single rewrite may not fully capture user intent.
  • This is useful at first-turn of question where single query rewrite does nothing.
  • The appropriate user would be paired with subset selection algorithm optimized for diversity, representativeness, etc.
4 Context Relevance
Granite 3.3 8b Instruct - Context Relevance is a LoRA adapter for granite-3.3-8b-instruct, that is fine-tuned for the context relevancy task:
Given (1) a document and (2) a multi-turn conversation between a user and an AI assistant, identify whether the document is relevant (including partially relevant) and useful to answering the last user question.
While this adapter is general-purpose and can even be used in cases where there is only one question, it is especially effective in RAG settings right after the retrieval model’s step, where the adapter can be used to identify documents or passages that may mislead or harm the downstream generator model’s response generation.
4.1 Intended Use
This LoRA adapter enables classification of relevant, partially relevant, and irrelevant documents for the final user query in a multi-turn conversation. The model has been trained to determine whether a document is useful and relevant for answering the last user question in a multi-turn conversation.
The classification output from the context relevancy model can be used in several downstream applications, including but not limited to:
  • Filter out irrelevant and misleading documents before sending them to a generator model in a RAG setting. By removing irrelevant documents upfront, the downstream generator model reduces the likelihood of outputting misleading responses. Moreover, removing irrelevant documents or passages can help reduce the total downstream context length, thereby improving inference time.
  • Signal to human annotators working in RAG settings which documents are irrelevant/relevant to the current turn of the conversation they are reviewing. Identifying such documents helps reduce the human annotator’s high cognitive load ( Hanafi et al. (2025)) involved in manually reading and reviewing several documents, especially in multi-turn conversations.
Model Input: The input to the model is a list of conversational turns and a list of documents, where each document is a dict containing the fields title and text. The turns in the conversation can alternate between the user and the assistant, and the last turn is assumed to be from the user. For every document in the list of documents, the model converts that document and the conversation into a string using the apply_chat_template function.
To prompt the LoRA adapter to determine context relevancy, a special context relevancy role is used to trigger this capability of the model. The role includes the keyword “context_relevance”: <|start_of_role|>context_relevance<|end_of_role|>
Model Output: When prompted with the above input, the model generates a JSON structure containing the context relevance output (“irrelevant”, “partially relevant”, “relevant”), as shown in the following example:
  "context_relevance": "irrelevant"
For an example describing how to use the intrinsic, refer to the documentation of the RAG agent library at https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib.
4.2 Evaluation
Our model was evaluated against several existing benchmarks to measure its ability to determine whether a document is relevant or not to the last user question in a multi-turn conversation.
4.2.1 Forming evaluation datasets for context relevancy
For each of the datasets, we created pairs of documents and questions and generated a label (based on the available labels in the dataset) as to whether the document is relevant or useful to answering the question.
For cases where all documents and question pairs are positive labels (the document is useful/relevant for answering the question), an instance with a negative label is generated by randomly pairing the document with an irrelevant question.
Given the dataset ground truth and the LoRA outputs, we then calculated the precision and recall scores for the relevant and irrelevant data instances.
4.2.2 Evaluation Datasets
We used queries and document sets from the following datasets for our evaluation:
  • Multi-turn conversational RAG benchmarks: MTRAG (Katsis et al. (2025)). The MT-RAG dataset comes with three settings of RAG retrieval: (1) Reference: perfect retriever (all passages are relevant to the question); (2) Reference + RAG: references passages + additional passages resulting in 5 total passages; (3) RAG: standard RAG setting (top 5 retrieved passages). We split the evaluation dataset according to the second and third settings and formed pairs of questions and passage sets. The dataset comes with pre-existing negative labels.
  • Question Answering datasets, that involve implicit-fact retrieval(Zhao et al. (2024)): CLAPNQ (Rosenthal et al. (2025)) and Drop (Dua et al. (2019)). The datasets contain pre-existing negative labels.
  • Several RAG benchmarks that have previously been used to evaluate Granite models: FinanceBench (Islam et al. (2023)), BioASQ [^4] We use the subset from https://huggingface.co/datasets/enelpol/rag-mini-bioasq (Thakur et al. (2021)), Open Australian Legal Corpus (Butler (2023)). We generated negative labels by random pairings.
  • Granite Guardian (Padhi et al. (2024)) datasets for context relevancy: HotpotQA and SquadV2. The evaluation datasets contained pre-existing negative labels, which were obtained via prompting a separate model over the original datasets.
4.2.3 Evaluation Results
We evaluated the context relevance LoRA adapter against meta-llama/Llama-3.3-70B-Instruct, ibm-granite/granite-guardian-3.1-5b, and the Granite model ibm-granite/granite-3.3-8b-instruct. For Llama, we prompted the model using the prompt proven to perform best for context relevancy( Rahmani et al. (2025)).
Evaluation Dataset Llama Granite Granite 3.3 Granite 3.3
3.3 70B Guardian Instruct LoRA
MTRAG (RAG) 0.93 0.90 0.93 0.96
MTRAG (Reference + RAG) 0.87 0.83 0.86 0.92
CLAPNQ 0.97 0.31 0.69 0.99
Drop 0.86 0.35 0.73 0.94
FinanceBench 0.92 0.44 0.66 0.99
BioASQ 0.98 0.83 0.94 0.99
Open Aus. Legal Corpus 0.93 0.08 0.60 1.00
HotpotQA + SquadV2 0.54 0.13 0.54 0.65
Table 10: Context relevance evaluation: Precision scores on the relevant labels
Evaluation Dataset Llama Granite Granite 3.3 Granite 3.3
3.3 70B Guardian Instruct LoRA
MTRAG (RAG) 0.41 0.21 0.29 0.73
MTRAG (Reference + RAG) 0.39 0.21 0.29 0.73
CLAPNQ 0.97 0.00 0.65 0.99
Drop 0.84 0.02 0.64 0.94
FinanceBench 0.89 0.00 0.34 0.99
BioASQ 0.87 0.00 0.53 0.99
Open Aus. Legal Corpus 0.92 0.00 0.36 1.00
HotpotQA + SquadV2 0.16 0.08 0.15 0.35
Table 11: Context relevance evaluation: Recall scores on the irrelevant labels
We report the precision scores on the relevant labels (Table  10) and the recall scores on the irrelevant labels (Table  11). We observe that the LoRA’s precision scores of the relevant labels improve on all the benchmarks compared to the precision scores on the Granite Instruct, Llama, and Granite Guardian models. We observe the same trend in the recall scores of the irrelevant labels for the LoRA adapter. Our evaluations show that the context relevance LoRA adapter is able to effectively identify relevant and irrelevant documents across several benchmarks.
4.3 Training Details
The training data was generated synthetically using a Mixtral model and documents from the CLAPNQ dataset (Rosenthal et al. (2025)). Based on this corpus, we generated several pairs of documents and multi-turn conversations along with a relevant/partially relevant/irrelevant label indicating the document’s relevance to the last user question. We then used Mixtral as an automatic judge to validate the generated labels and filter out noisy samples.
The LoRA adapter was fine-tuned using PEFT under the following regime: rank = 32, learning rate = 5e-6, number of epochs = 25.
5 Answerability Determination
Granite 3.3 8b Instruct - Answerability Determination is a LoRA adapter for ibm-granite/granite-3.3-8b-instruct fine-tuned for binary answerability classification task. The model takes as input a multi-turn conversation and a set of documents, and classifies whether the user’s final query is answerable or unanswerable based on the available information in the set of input documents.
5.1 Intended Use
For an example describing how to use the intrinsic, refer to the documentation of the RAG agent library at https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib.
5.2 Evaluation
5.2.1 Answerability Classification
    BigBird w/ MLP 49.2 68.5 57.3 48.0 29.2 36.3 46.8
    LLaMA 2-7B 72.2 71.0 71.6 71.4 72.6 72.0 71.8
    Granite 3.3-8b LoRA 88.1 59.3 70.9 69.3 92 79 75
BigBird w/ MLP 69.6 77.6 73.4 70.1 60.8 65.2 69.6
LLaMA 2-7B 86.9 89.4 88.2 87.3 84.5 85.9 87.1
Granite 3.3-8b LoRA 89.8 91.8 90.8 90.3 87.9 89.1 90
5.2.2 Comparing LoRA Adapter vs. Vanilla Granite for Answer Quality
We compare the performance of Granite 3.3-8b Instruct vs. Granite 3.3-8b LoRA adapter on a subset of MT-RAG Benchmark in Table 14. In this setup, each query is paired with only 5 retrieved passages as context. The true answerability label for each query indicates whether the query is answerable with respect to the retrieved context.
  • Joint Answerability-Faithfulness Score (JAFS):
    The LoRA adapter achieves a 17% lift on this metric - rewarding the model for correctly abstaining on unanswerable queries and for being faithful when it chooses to answer.
Model Unans. F1 Ans. F1 Unans. Recall Ans. Recall JAFS
Granite 3.3-8b Instruct 17 77 10 99 49
Granite 3.3-8b LoRA 69 81 67 82 66
Table 14: Comparison of Granite 3.3-8B Instruct vs. LoRA Adapter on Answerability and Faithfulness metrics using MT-RAG Benchmark.
5.3 Training Details
6 Passage Reranking
Granite 3.3 Instruct - Passage Reranking is a prompt-based intrinsic for reranking retrieved passages. It takes the output of the retrieval step as input and returns a reranked (subset of the) retrieved passages which can be then used for generation.
6.1 INTENDED USE
The reranking is done in a tournament style by prompting an LLM to compare paired passages and return the label of the preferred one. Users can provide their own prompt; otherwise the following default prompt is used:
1You are a smart and helpful AI assistant with in-depth knowledge about how people search for information using search engines. In this task, you are given two passages, and a query, your job is to judge which passage is relatively more suitable to answer that query. The first passage will start with "passage A" and the second passage with start with "passage B". Output the preferred passage index, i.e. A or B and followed by an explanation, if none of the passage answer the query directly, pick the one that has more relevant information.
When the number of reranking passages is less than 10, the intrinsic compares all possible pairs and ranks by each passage’s win count. All or multiple pairs are grouped into single batch for efficiency. When the number of reranking passages is larger than 10, the intrinsic forms pairs using the original rankings and only advances the winning passage from each pair to the next round. If there are an odd number of documents, the last document is dropped for simplicity.
For an example describing how to use the intrinsic, refer to the relevant documentation in Granite IO at https://github.com/ibm-granite/granite-io/tree/main/notebooks.
6.2 EVALUATION
We provide the re-ranking result for the popular BEIR (Thakur et al. (2021)) benchmark as reference. By closely examining the re-ranking results, we find that many of the re-ranked top passages across different tasks are at least as good as the gold passages, although they are not counted as gold. We suspect this contributes to the performance drop in some tasks.
NDCG@10 granite-embedding-30M Granite 3.3 Instruct (rerank top 40)
trec-covid 0.631 0.724
nq 0.516 0.586
fiqa 0.369 0.371
scifact 0.713 0.742
msmarco 0.307 0.322
hotpotqa 0.629 0.652
nfcorpus 0.337 0.367
touche2020 0.240 0.223
scidoc 0.225 0.210
dbpedia 0.360 0.405
fever 0.855 0.848
climatefever 0.303 0.301
CQADupstaceRetrieval 0.443 0.459
quora 0.867 0.805
arguana 0.567 0.347
Table 15: NDCG@10 scores on the retrieval tasks in the BEIR benchmark, prompting Granite 3.3 Instruct.
Unlike most generation tasks, reranking only needs to ouput 1 single preferance token, A or B. Therefore, the processing time is significantly less than regular generation and it only adds a small overhead in the end to end RAG pipeline. Possible future work includes support for reranking a set of passages rather than only a pair.